OpenStudy (anonymous):

A scale drawing of a map shows that 2 inches = 24 miles. How many miles are represented by 1/4 inch on the map

OpenStudy (omarbirjas):

1/4=.25

OpenStudy (omarbirjas):

and for every 2 inches, you have 24 miles

OpenStudy (omarbirjas):

we can write it like this 2/24 x .25/n

OpenStudy (omarbirjas):

n is what we're trying to find

OpenStudy (omarbirjas):

all we have to do is cross multiple. 2xn=24x.25

OpenStudy (omarbirjas):

2n=6

OpenStudy (omarbirjas):

now we have to get rid of the 2 so what we do is devide both sides by 2 and what do we get

OpenStudy (omarbirjas):

2n/2=n and 6/2=3

OpenStudy (omarbirjas):

n=3